A hypothesis test is conducted to test whether or not the population correlation is zero. If the sample consists of 38 observations and the sample correlation coefficient is -0.86, what is the computed value of the test statistic? Round to two decimal places.

Question 16 options:

1)

-1.96

2)

-10.12

3)

10.12

4)

2.00
Answer #1

Solution :

Give that,

r = -0.86

n = 38

degrees of freedom = n - 2 = 38 - 2 = 36

Test statistic = t

= TV (n − 2)/ \(1 – 12)

= -0.86 * (\sqrt36) / \sqrt (1 - (-0.86)2)

= -10.12

Test statistic = -10.12

Option 2)